TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.BorisBellesia on 2006-12-12 - 20:28

60A converter tests - Tuesday 12th Dec.

Advancement Tue. 12 Dec.:

  • AB-PO activities ended in the whole arc
  • Polarity tests ended in the whole arc

Activities Wed. 13th Dec.:

  • TS-EL pre-conneciton in C32-34_L8 / C 12-34_R7; no problems for coactivities
  • if it is possible: El_QA.doc in C18-28 _L8; co activities limited - a trolley is used

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-12 - 15:21

RAT for Point 5 in CCC, Dec 12

PDF RAT_20061212.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x12x15x08/RAT_20061212.pdf

PDF UJ56-RR57_IR-Report_2006-12-12_24H-Burnin.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x12x15x08/UJ56-RR57_IR-Report_2006-12-12_24H-Burnin.pdf

Comments: NEXT MEETING WILL BE HELD ON DEC,18 AT POINT 5

Please, find also included the IR analysis rapport, kindly provided by Jeff Thomsen.

Open Issues

  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)
  • 06.12 Permanent labels for RQ converters in RR57 (according to ECR)
  • 07.12 Access key to electrical distribution room in UJ56

Heat Runs and Test Schedule

  • 01.12 Squeeze test DONE (all the power converters involved have been tested)
  • 06.12 8-h heat run DONE
  • 07.12 Fire detection test DONE
  • 08.12 Full power ventilation tests in RR57 – from 8:30 till 12:00 DONE
  • 08.12 Afternoon: Polarity Test_1st part TO BE RE-SCHEDULED
  • 11.12 10am: 24-h heat run DONE
  • 13.12 9.00: UPS test and Electric cut (AUG simulation).
  • 13.12 Polarity Test_2nd part TO BE RE-SCHEDULED
  • 14.12 13kA energy extraction test
  • 15.12 14.00: New sequencer test

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-12 - 12:40

Preparation for Cool Down RAT sec78 - Dec 12

PDF RAT_20061212.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x12x12x38/RAT_20061212.pdf

Comments:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum
  • 12.12.06 One Pressure Gauge in DFBs cannot be read (F.Millet)
Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 08.12.06 DFBXG 100V test on current leads (S.Feher)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

MircoCoccoli - 12 Dec 2006 - 12:38

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-12-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login